OpenStudy (anonymous):

If f(x + 3) = \(x^2 + 2x + 4\), find f(x). Do I just plug in x + 3 for x?

OpenStudy (anonymous):

Replace x + 3 by x.. So put x - 3..

OpenStudy (anonymous):

Wait, why would that work? I just want to understand the dynamics of why it works.

OpenStudy (anonymous):

put x - 3 inplace of x because f((x-3)+3)=f(x-3+3)=f(x)

OpenStudy (anonymous):

If you replace there x by x -3: Then: \[f(\color{green}{x} + 3) = f(\color{blue}{x - 3} + 3) = f(x)\]

OpenStudy (anonymous):

Oh I get it now... f(x) = x^2 - 4x + 7?

OpenStudy (anonymous):

Well Done..

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

Hey right this time you became 93.. Congratulations..

OpenStudy (anonymous):

Thank you again :)